OpenStudy (anonymous):

find the amount of time in years that it will take an investment of $1000 to double at 8%interest compounded continuosly. (hint A=Pe^rt

OpenStudy (anonymous):

Hi @CesarLovesNadia - so you know the formula you need to use, correct? Are you able to put them together?

OpenStudy (anonymous):

no im not thats why its here i need help

OpenStudy (anonymous):

No problem, so the interest formula for compound interest is what you have listed. So, A is the number we are looking for, we have the rest. So, A = P (principal amount) x e (Napier's #, roughly 2.7183) ^ r (interest rate) x t (number of years). So, how about you try and put in the numbers you know, and I can help you with anything you don't understand?

OpenStudy (anonymous):

For the 'e' - your calculator will actually have this number built in, so don't worry about it for now :)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

how many years would it be?

OpenStudy (anonymous):

@CesarLovesNadia - You have to learn how to do these problems - I'm not here to give you the answer. I WILL help you get there though, I promise. Can you try and put the variables from the word problem into the equation now?

OpenStudy (anonymous):

i diid i put 1000*2.7183^(0.08*1)

OpenStudy (anonymous):

it came out to be like 1000 something

OpenStudy (anonymous):

No worries - and good start. The problem states that you need to find the time, t, for your money to double. Therefore, you know that A = 2xP = 2000. So, your equation looks like this: \[2000 = 1000*e^(.08*t)\]

OpenStudy (anonymous):

so then wait

OpenStudy (anonymous):

the answer would be 2000 then?

Parth (parthkohli):

Colm, it'd be \(2000 = 1000 \times e^{(0.8 \times t)}\) Sorry for interfering, guys.

OpenStudy (anonymous):

thanks

Parth (parthkohli):

@CesarLovesNadia You should appreciate Colm for helping you :) P.S. He's the major part behind OpenStudy ;)

OpenStudy (anonymous):

Heh, that's what I was writing - guess it showed up wrong. And no, @CesarLovesNadia you need to find the TIME, t in this equation, right? "find the amount of time in years"

OpenStudy (anonymous):

but what is it then?

OpenStudy (anonymous):

just help me and give me the answer

OpenStudy (anonymous):

@CesarLovesNadia - sorry, but you should read the Code of Conduct ( http://openstudy.com/code-of-conduct) - we don't just give the answer. I can help you understand it though :) So now, you have to solve for T in that equation I and Parth just gave you. Can you start that?

OpenStudy (anonymous):

so would it be 2?

OpenStudy (anonymous):

is it 2?

Parth (parthkohli):

Heh—don't give me the credit. Keep it on, guys!

OpenStudy (anonymous):

help me with another one

OpenStudy (anonymous):

Given your equation 2000 = 1000 x e^(.08xT) We divide both sides by 1000 to give us this: 2 = e^(.08xT)

Parth (parthkohli):

Wolfram time! http://www.wolframalpha.com/input/?i=e%5E%280.8t%29+%3D+2

Parth (parthkohli):

Your Math is pretty good, Colm!

OpenStudy (anonymous):

@CesarLovesNadia - you'll probably want to use that link to get the real answer. Hint - it's not two.